OpenStudy (thomas5267):

Find the remainder when \(67^{101}\) is divided by 65.

ganeshie8 (ganeshie8):

Hint \[67\equiv 2\pmod{65}\]

OpenStudy (thomas5267):

I thought 65 was a prime number, applied Fermat's little theorem and failed miserably...

ganeshie8 (ganeshie8):

Thats typical but you may apply euler's theorem for composite moduli

ganeshie8 (ganeshie8):

\(\phi(65)=\phi(5\times 13)=(5-1)(13-1)=48,\)therefore \[a^{48}\equiv 1\pmod{65}\] for all \(a\) relatively prime to \(65\)